OpenStudy (anonymous):

Please, someone explain this to me; Solve for R in 1 equation (in terms of S and T and k): R is inversely proportional to S and R is directly proportional to the cube of T

OpenStudy (bahrom7893):

r = t^3/s

OpenStudy (anonymous):

in terms of t, s, AND k

OpenStudy (bahrom7893):

actually i think it's: r = k*t^3/s

OpenStudy (anonymous):

"Please, someone explain this to me;"

OpenStudy (bahrom7893):

directly proportional means equal to pretty much.. So if a is directly proportional to b, then a=b

OpenStudy (bahrom7893):

inversely proportional means equal to 1/ that number

OpenStudy (anonymous):

r = k/s r = kt^3 I got this far but I'm not sure from here

OpenStudy (bahrom7893):

No u just keep one constant

OpenStudy (anonymous):

You are incorrect about inverse and direct direct: y = kx inverse: y = k/x

OpenStudy (bahrom7893):

It's proportional to one and the inverse of the other at the same time.. I'm not incorrect, ignore the constant for now

OpenStudy (anonymous):

why?

OpenStudy (bahrom7893):

The constant is ALWAYS put all the way at the end

OpenStudy (anonymous):

How come?

OpenStudy (bahrom7893):

First you finish your equation, then you stick it on top.. unless the constant is initially given, because then when u multiply the values, u'll end up with k^2, which is still a constant, so it's still k

OpenStudy (bahrom7893):

I mean just put it out in front all the way at the end

OpenStudy (bahrom7893):

amistre i dunno how to explain this...

OpenStudy (anonymous):

Oh, that makes sense, if you put it on both, then you would have 2 ks, is that what you are saying?

OpenStudy (amistre64):

directly porportional is when the value of one increases the other decrease: x*y = 5 is directly porportional; if x gets big, y has to shtink to retain the same value

OpenStudy (bahrom7893):

amistre probably has an explanation lol

OpenStudy (bahrom7893):

wait really???

OpenStudy (anonymous):

i didnt learn it that way

OpenStudy (anonymous):

i learned direct both increased and inverse one increased as the other decreased

OpenStudy (bahrom7893):

amistre directly proportional is x=y, ur equation would give x = 1/y

OpenStudy (amistre64):

indirectly is when the values increase or decrease at the same time y/x = 5 is an indirect porportion since an increase in one results in the increase of the other to retain the same value

OpenStudy (amistre64):

i could be hazy on it, so let me recheck the google :)

OpenStudy (bahrom7893):

inversely proportional is x*y=constant

OpenStudy (anonymous):

amistre i think you got it backwards

OpenStudy (amistre64):

its possible :) but if so then youve got a pretty good bead on it

OpenStudy (bahrom7893):

lol yea.. he prolly did this years ago

OpenStudy (anonymous):

I just don't understand how to get one equation out of the two

OpenStudy (amistre64):

just relate them to it in the same manner

OpenStudy (anonymous):

will you show me?

OpenStudy (bahrom7893):

All right listen: R = kT^3/S

OpenStudy (amistre64):

What does: R is inversely proportional to S imply then? since your right lol

OpenStudy (anonymous):

r = k/s

OpenStudy (bahrom7893):

Inversely is 1/S

OpenStudy (bahrom7893):

or k/s

OpenStudy (amistre64):

good; and what does and R is directly proportional to the cube of T imply?

OpenStudy (anonymous):

r = t^3

OpenStudy (anonymous):

kt^3 sorry

OpenStudy (amistre64):

then we put those toghter to form: \[r=k\frac{t^3}{s}\]

OpenStudy (anonymous):

how exactly?

OpenStudy (amistre64):

we just went over the how part

OpenStudy (amistre64):

r is related to both s and t in these matters

OpenStudy (amistre64):

manner

OpenStudy (amistre64):

are you related to your mother and father?

OpenStudy (anonymous):

but kt^3 = k/s kst^3 = k st^3 = r

OpenStudy (anonymous):

Yes, I am. Hopefully

OpenStudy (amistre64):

we are not equating k with k; we are relating how r is related to t and s

OpenStudy (amistre64):

k is just a byproduct that makes the whole thing consistent

OpenStudy (amistre64):

mother = father is what your trying to determine as an analogy :)

OpenStudy (anonymous):

?

OpenStudy (amistre64):

stop equating the 2 parts and unite them under the constant of parent

OpenStudy (anonymous):

I'm sorry

OpenStudy (anonymous):

is mom = dad what I was doing, or what i need to be doing?

OpenStudy (amistre64):

what you were doing, so stop it lol

OpenStudy (anonymous):

okay, what am I trying to do?

OpenStudy (amistre64):

r is related to s how? r is related to t how?

OpenStudy (anonymous):

r = k/s r = kt^3

OpenStudy (amistre64):

not "k"; just s and t "k" is not a relation to r, its a consequence of it

OpenStudy (anonymous):

but isnt r a consequence of s and k?

OpenStudy (amistre64):

as r changes, k stays the same so it of no relation as r changes both s and t change, they are related to r

OpenStudy (anonymous):

Ah! Well worded! k always stays the same because it's a constant, but as one variable changes the other has to since k cant change!

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

But now what do i do?

OpenStudy (amistre64):

since r is related to s and t in some manner, combine the terms for s and t into their proper relations with r when mom and dad are in the same room, their relationship with you doesnt change does it?

OpenStudy (amistre64):

aunt betty comes over for dinner does the family disintegrate?

OpenStudy (anonymous):

no it does not.

OpenStudy (anonymous):

does aunt betty = k?

OpenStudy (amistre64):

then r keeps its same reltaion to s and t as it always has lol, aunt betty = some other relation to r that is not mentioned k is just the house, or dinig room, or whatever keeps the family in the same place packed together

OpenStudy (amistre64):

r = k(relations)

OpenStudy (anonymous):

if r doesnt = k/s then how would you write an equation for it, just replace it with a random constant? r = 1/s? then r = 1t^3?

OpenStudy (amistre64):

r = 1/s is a stated relation; but we all know that 1/5 so we have a constant of variation that always allows the relation to stay solid

OpenStudy (amistre64):

soo slow on here

OpenStudy (anonymous):

I'm sorry I just don't understand, thanks for taking your time though. where'd you get 1/5?

OpenStudy (amistre64):

does 3 = 1/5? was what i typed but got lost in the etheral void of openstudy

OpenStudy (amistre64):

say when r = 3, s=5 what makes this relation solid?

OpenStudy (anonymous):

no, so how do we write an equation for r in terms of s that excludes k?

OpenStudy (amistre64):

the final result has a k so that the relations are consistent.

OpenStudy (amistre64):

3 = 15/5 20 = 15/something else lol

OpenStudy (anonymous):

well, if the constant of variation is given, you cant change the other two variables to not match up with it, they have to follow r=k/s, if you have 10 = 20/s, s has to be 2, if we have 10 = k/10, k has to be 100

OpenStudy (amistre64):

the relations dont change; so pack everyone together in the same k(..) to keep it mathematically consistent and happy

OpenStudy (amistre64):

10 to 10 isnt the proper relation then and its a fraud if you have to change the k

OpenStudy (anonymous):

you cant change k though, its a constant

OpenStudy (anonymous):

you have to change the variables to match up with it

OpenStudy (amistre64):

right, so 10 to 10 is not a valid relation between r and s in that case since you had to change k to match it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how do you combine the relations? and why just randomly multiply k out in front?

OpenStudy (amistre64):

when we have values for r and s and t then we can determine what the proper k will be, but until then we just use it as a genereal fill in

OpenStudy (anonymous):

how do you know that k * combined relations = r?

OpenStudy (amistre64):

it doesnt say that r is related to s only when t is not around; and when t is around s leaves so that r is related to t and such; it says that r is related to s and t together

OpenStudy (amistre64):

in this manner as such and such

OpenStudy (anonymous):

okay so I get the r = t^3/s, but how do you know where to put k?

OpenStudy (amistre64):

k is just the constant that ties it all together so that no random false relations stick their head in at the family reunion it goes where its always gone; as a multiplier to keep the pease

OpenStudy (anonymous):

It was not multiplied in r = k/s

OpenStudy (anonymous):

Just forget about it. I don't think I'll get this. Thanks for your help and time.

OpenStudy (anonymous):

I'll just teach myself by saying that r = k/s and r = kt^3 then wanting t somewhere in r = k/s, so asking myself what relation does k have with t, and saying it's multiplied by t^3 so r = kt^3/s, similarly, ks relation with s is k is divided by s so in r = kt^3 becomes r = kt^3/s. Not sure if that is the way to do it but I like that way.